OpenStudy (anonymous):

(b+9) x (b+3) explain/ check if my answer is right

OpenStudy (anonymous):

so do i only do 9x3?

OpenStudy (anonymous):

which is 27

OpenStudy (anonymous):

no use this formula (a+b)(a+b)= a^2+2ab+b^2

OpenStudy (mayankdevnani):

no !! there is also a variable i.e b you have to multiply this variable also

OpenStudy (anonymous):

FOIL method

OpenStudy (anonymous):

so it would be b^2+12b+27

OpenStudy (mayankdevnani):

\[\large \bf (b+9)(b+3)=b(b+3)+9(b+3)\] \[\large \bf b^2+3b+9b+27\] \[\large \bf b^2+12b+27\]

OpenStudy (anonymous):

oh okay i think i get it now O.o

OpenStudy (mayankdevnani):

good :)

OpenStudy (anonymous):

so for (x+2) x (x-3) i would use the same formula @mayankdevnani

OpenStudy (mayankdevnani):

yeah !!!

OpenStudy (mayankdevnani):

REMEMBER :- \[\large \bf (a+b)(c+d)=a(c+d)+b(c+d)\]

OpenStudy (mayankdevnani):

@mcnabkt

OpenStudy (anonymous):

so then its x^2+ 2b-3 ugh my brain hurts like idk if i have it right so far

OpenStudy (anonymous):

@gorv

OpenStudy (anonymous):

i just get confused with the a and b letters like idk

OpenStudy (mayankdevnani):

a,b,c,d are numbers

OpenStudy (anonymous):

well if there are only two numbers then whats c?

OpenStudy (anonymous):

and d?

OpenStudy (mayankdevnani):

c and d are variables or numbers

OpenStudy (mayankdevnani):

and also a and b are variable or numbers

OpenStudy (mayankdevnani):

Like in your question :-